(15 Q3. nts) 13 3 X 7 → 3 32 Determine the x- and y-coordinates of the centroid of the shaded area. Show your calculat
Answer #1

Page no 101 13 26 GA 62 14 .93 31 We will divide the given Area 1311 into 4, areas. 25 GI 14 G4 (63 34 6 . 4r 11 8 4 X- coor(fage no 2 where A = Area of Big rectangle 13 x 14 - 182 inch a = location of centroid of rectangle from Yaxis 1, 외에 13 2 6.5(Page no Y co-ordinate of controid. Y TA:y: TAI Any + A22 A393 - A4.94 Asa A2 A3- Ay where 182 inch Y = Location of controidPagenoy the equation Substituting all values in J = (182*7) + (50*4.07) - (4.137 x 1.273)-(12x5) 182 + 56 14.137 - 12 9 = 6 8In this numerical,we want to find the x and y co-ordinate of the shaded area.So we will divide the figure into 4 areas.Then we will use the formula of centroid of composite area and find the required coordinates.

I have written the page number for your reference.

If you like my answer then please give me a " Thumbs Up ".It motivates me to solve more questions and help more number of students.

Thank You.